and E The accumulated depreciation at the end of year 7n$
Answer #1
Req 1)
Formula for calculating revised depreciation expense:
(Original Cost of asset - Accumulated Depreciation at the end of 6th year) / Estimated remaining useful life of asset
[32000 - (32000 * 6 / 16)] / 5    = (32000 - 12000) / 5 = 4000 per year
Journal entry to record 7th year depreciation on the equipment:
Date Accounts and explanation Debit $ Credit $
Dec 31 7th year Depreciation expense 4000
Accumulated Depreciation 4000
(being depreciation expense booked on the revised rate )
Req 2)
Accumulated Depreciation at the end of 7th year =
depreciation for first 6 years + Revised depreciation for 7th year =
(2000 * 6) + 4000 = $16000